# Variational Mode-Driven Graph Convolutional Network for Spatiotemporal Traffic Forecasting

### v1

@@ Abstract
This paper focuses on spatio-temporal (ST) traffic prediction traffic using graph neural networks. Given that ST data consists of non-stationary and complex time events, interpreting and predicting such trends is comparatively complicated. Representation of ST data in modes helps us infer behavior and assess the impact of noise on prediction applications. We propose a framework that decomposes ST data into modes using the variational mode decomposition (VMD) method, which is then fed into the neural network for forecasting future states. This hybrid approach is known as a variational mode graph convolutional network (VMGCN). Instead of exhaustively searching for the number of modes, they are determined using the reconstruction loss from the real-time application data. We also study the significance of each mode and the impact of bandwidth constraints on different horizon predictions in traffic flow data. We evaluate the performance of our proposed network on the LargeST dataset for both short and long-term predictions. Our framework yields better results compared to state-of-the-art methods.
Index Terms:
Traffic data, Spatio-Temporal, Forecasting, Graph Neural Network, Decomposition
I Introduction
@@
Traffic Management Systems (TMS) play a vital role in the growth of urbanization. The increasing number of vehicles on the road consistently pressures traffic planning and management. An Intelligent Transportation System (ITS) enhances the efficiency, reliability, and safety of transportation systems [1]. Accurate predictions help avoid congestion and improve traffic infrastructure, mitigating issues like air pollution, fuel and time wastage, and accidents. These problems directly impacts on the human health, daily life, social life, and psychological states, including lack of control, fatigue, time pressure, and stress. For effective online monitoring and prediction, smart cameras or sensors must be installed at crowded intersections, and path-planning user applications should guide individuals away from traffic jams. The traffic counts or flow holds statistical significance in data analytics, allowing predictions of increasing vehicle numbers over time and planning for traffic infrastructure accordingly. Socioeconomic indicators and policies also play roles in traffic management. The policies based on speed or flow can be implemented to prevent road crashes, motivating us to address the problem of developing accurate and robust traffic prediction methods. In this paper, we are addressing traffic prediction using a deep learning method.
I-A Related Work
@@
In traffic forecasting, both parametric and non-parametric models are fundamental approaches for predicting future states. The choice of model depends on the size and complexity of the data, as well as the available computational resources. Parametric models include time series analysis, parametric regression, and Kalman filter models. Time series analysis models, such as the auto-regressive integrated moving average (ARIMA), seasonal ARIMA (SARIMA), and exponential smoothing. The non-parametric approaches including K-nearest neighbors, decision trees, and Support Vector machines (SVM) have been widely used.
@@
The Box-Jenkins ARIMA model, which transforms non-stationary data to stationary data using the differences in the time series, is frequently incorporated in traffic volume prediction [2]. Another approach for prediction in [3], uses measured state data to forecast traffic flow, minimizing prediction errors based on the assumption of data stationarity (Kalman filtering theory). These parametric methods are interpretable but struggle to model complex and non-stationary traffic data effectively. Non-parametric models utilize distance metrics to determine the nearest neighbor of the current state to the historical observation in feature space [4]. These methods are rich in data requirements, less interpretable, and computationally complex. Both parametric and non-parametric methods utilize historical observed data but do not incorporate spatial information.
Due to recent advancements in deep learning, significant improvements have been made in traffic forecasting applications. The state-of-the-art methods in traffic prediction consider various factors such as spatial topology construction, spatial dependency, temporal dependency, and external factors. These advanced methods are designed to better handle the complexities of modern traffic data and provide more accurate predictions over both short (refer to horizons of up to one hour) and long term (beyond one hour) [5].
Spatial topology construction
@@
The non-Euclidean geodesic information of the sensor network is incorporated in the model using graph neural networks (GNNs). If the graph topology is defined based on the distance between the nodes, this representation is known as a static graph. In an adaptive graph, the structure of the graph is constructed based on the temporal evolution of the data. For spatial graph modeling from the dynamic feature, [6][7][8] determines a self-adaptive adjacency matrix using node embeddings end-to-end learning mechanism. Dynamic graph convolutional recurrent network [9] incorporates the weighted combination of both static and dynamic graph convolution. The semantic adjacency matrix is computed using dynamic time warping distance by incorporating the time series of nodes [10]. The binary adjacency matrix represents the presence of the connection between nodes [11]. The road network is also included in distance-based graph structure [12].
Spatial dependency
@@
To extract the spatial feature of traffic flows from the road network matrix, trafficGAN [13] implements deformable convolution [14]. Deformable convolution is used to model geometric variations. To capture the spatial dependency, mostly CNN [15][16] and GNN  [12][17][7] are employed in prediction applications. The self-attention mechanism [18] is also used in the modeling of spatial dependency. In the most recent literature on traffic prediction [19][11], the combination of attention and GNN is also encouraged to determine the spatial correlation. The meta-graph attention which is based on graph attention (GAT) is applied to determine the spatial correlations from traffic data  [20]. Chebyshev spectral convolution (ChebNet) is incorporated to learn spatial dependencies for spectral graph convolution [21]. To overcome the smoothing issues in GNNs in deeper networks that impact the long-range dependencies, the spatial-temporal graph ordinary differential equation (STGODE) block is proposed in [10]. STGODE captures the spatial and temporal dynamic correlations based on ODE.
@@
Traffic prediction involves time series data used to capture sequential relationships. RNN and their optimized variants, such as LSTM and GRU, are commonly employed. These models are capable of capturing the long-range temporal dependencies [9]. However, combining of dynamic graphs with RNNs make the model computational expensive due to a large number of trainable parameters. Temporal convolution filters are also used to extract the temporal information from the data. The attention mechanism [22] in the spatio-temporal (ST) block is commonly used in predictive models [11][19]. A combination of RNNs along with the attention mechanism is also employed. GRUs handle the short-term dependencies, while the self-attention mechanism models the long-term dependencies [8]. Three different types of attention– temporal, spatial, and channel– are applied to capture spatio-temproal correlations [18]. However, this proposed mechanism increases the computational, time complexity, and number of parameters compared to CNNs. CNNs and TCNs have proven more efficient in capturing complex spatio-temporal relationships. The transformer encoder-decoder architecture is implemented for the temporal pattern in [17][20]. Typically, the encoder is implemented to learn historical traffic information and the decoder is incorporated to make predictions.
@@
Traffic data consists of complex dynamics that depend on various factors socio-economic, land uses land covers, population density, road capacity, and cultural or public events. To apply self-attention to the point of interest (POI) of each region, POI-MetaBlock is included in deep neural network (DNN) [23]. To effectively capture temporal correlation, the external weather data is incorporated in DNNs for traffic flow forecasting in [24]. The non-traffic data such as rainfall [25] is also included in the model. Information such as holidays, weather, and nearby POIs are served as the input to the Large Language Models (LLMs) [26]. These factors serve as the features of DNNs that assist the network in extracting the spatial and temporal correlation from the traffic data.
@@
Fourier domain analysis is also used in traffic forecasting, where the traffic data is segregated into frequency components. However, the limitation of the Fourier Transform is assumed that the data should be stationary. So, many signal processing-based methods have been introduced for the decoupling of signals. Various methods are commonly used for signal decomposition, including the well-known methods Empirical Mode Decomposition (EMD) [27], Ensemble EMD (EEMD) [28], Complete Ensemble EMD with Adaptive Noise (CEEMDAN) [29], Multivariate EMD (MEMD) [30], and Variational Mode Decomposition (VMD) [31]. The traffic data is decomposed into periodic, residual, and volatile components using the EMD method. A Bi-directional LSTM model is incorporated to model the volatile component [32]. A hybrid model, combining Improved CCEMDAN (ICCEMDAN) and GRU, is proposed to predict the parking demand [33]. Short-term traffic flow is predicted using VMD and LSTM [34]. The encoder-decoder architecture is incorporated for feature extraction from mode functions decomposed by VMD in short-term traffic flow prediction [35].
I-B Contribution
The state-of-the-art methods focus on short-term prediction using the decomposition methods. The impact assessment of noise and the effectiveness of each mode need to be considered in the prediction analysis using DNNs. In this work, we study both short-term and long-term prediction using a hybrid model (VMD + ASTGCN) approach. The main contributions of this work are summarized below,
  • In our hybrid framework for spatio-temporal forecasting, we propose to decompose the spatio-temporal traffic data into modes representing the granularity of the data using VMD. The spatial and temporal attention are applied to the decomposed features, followed by Cheby-Net graph convolution and time convolution to predict the future states over the long term.
  • We conduct an in-depth analysis of our proposed architecture, and formulate a method to determine the best-fit value of the number of modes in the real-time dataset. Additionally, we highlight the impact assessment of different hyper-parameters on the long-term prediction performance of the model.
  • We evaluate the performance of the proposed architecture on the LargeST[36] dataset across three different regions using three metrics mean absolute error (MAE), root mean square error (RMSE), and mean average percentage error (MAPE). The collective nodes of the LargeST dataset are $6902$, sampled at $15$ minute time interval over one year (containing $241$ million spatio-temporal observations). As an illustration, Fig. 1 demonstrates the prediction loss of our proposed model on the SD region.
The rest of the paper is organized as follows. Section II describes the adopted notation and formulates the problem statement. Section III presents the detailed architecture of the network used in our framework along with the mathematical formulation of the decomposition method. Section IV presents the experimental analysis and the improvements enabled by our methods before Section V concludes the paper.
@@ I-B Contribution
Figure 1: This figure contains the visualization that shows the temporal normalized loss variations on SD region for (a) Horizon 3 (b) Horizon 6 (c) Horizon 12. (For proper visualization of animation, please open this document in Adobe Reader).
II Preliminaries
II-A Road Network
@@
In graph theory, a road is modeled by considering different nodes at distinct joints, vertices, or intersections. Each node of the road is connected to another segment or node. The interconnectivity and distance among nodes are incorporated in a graph referred to as a road network, defined as $\mathcal{G}=(V,E,A)$, where $V$ is a set of entities or nodes, $E$ represents the connection between these entities, and $A\in\mathbb{R}^{N\times N}$ is the weighted adjacency matrix. Here, $N$ represents the total number of nodes in a traffic prediction application, which could be either the number of sensors or vertices of a road network.
II-B Temporal Traffic Network Graph
@@
Traffic flow, the number of vehicles per unit of time in a road segment, is a key measure of traffic congestion, which also depends on the width or number of road lanes. A temporal dynamic graph is an arrangement of graph snapshots [$\mathcal{G}^{(1)},\mathcal{G}^{(2)},\dots,\mathcal{G}^{(T_{w})}$] where $\mathcal{G}^{(t)}=(V^{(t)},A^{(t)},X^{(t)})$ includes vertices $V^{(t)}$, an adjacency matrix $A^{(t)}$, and a feature matrix $X^{(t)}$ [37]. Here, $T_{w}$ is the time window depending on the horizon of the signal. The feature matrix $X^{(t)}$ may include traffic information and road network characteristics. This structure, known as the temporal traffic network graph (TTNG), explains the traffic flow among different nodes from time $[1,T_{w}]$. As an illustration, Fig. 2 depicts the sequence of temporal graphs for a traffic network.
@@
Figure 2: Left side illustrates the spatial distribution of nodes in the San Diego region, as represented in the LargeST dataset. On the right side, the TTNG structure is depicted, where the historical and current input graphs are sampled at time intervals $t=\{1,2,\dots,T_{w}\}$. The output graphs are then predicted for $t=\{T_{w}+1,T_{w}+2,\dots,T_{w}+N_{H}\}$.
II-C Problem Formulation

### v3

@@ Abstract.
This paper focuses on spatiotemporal (ST) traffic prediction using graph neural networks (GNNs). Given that ST data comprises non-stationary and complex temporal patterns, interpreting and predicting such trends is inherently challenging. Representing ST data in decomposed modes helps infer underlying behavior and assess the impact of noise on predictive performance. We propose a framework that decomposes ST data into interpretable modes using variational mode decomposition (VMD) and processes them through a neural network for future state forecasting. Unlike existing graph-based traffic forecasters that operate directly on raw or aggregated time series, the proposed hybrid approach, termed the Variational Mode Graph Convolutional Network (VMGCN), first decomposes non-stationary signals into interpretable variational modes by determining the optimal mode count via reconstruction-loss minimization and then learns both intramode and cross-mode spatiotemporal dependencies through a novel attention-augmented GCN. Additionally, we analyze the significance of each mode and the effect of bandwidth constraints on multi-horizon traffic flow predictions. The proposed two-stage design yields significant accuracy gains while providing frequency-level interpretability with demonstrated superior performance on the LargeST dataset for both short-term and long-term forecasting tasks. The implementation is publicly available on GitHub.
spatiotemporal prediction, traffic forecasting, graph neural networks, variational mode decomposition, deep learning, time series analysis
journal: TISTccs: Applied computing Transportationccs: Information systems Spatial-temporal systemsccs: Computing methodologies Neural networks
1. Introduction
@@
Traffic Management Systems (TMS) play a vital role in the growth of urbanization. The increasing number of vehicles on the road consistently pressures traffic planning and management. An Intelligent Transportation System (ITS) enhances the efficiency, reliability, and safety of transportation systems (Haydari and Yılmaz, 2020). Accurate predictions help avoid congestion and improve traffic infrastructure, mitigating issues like air pollution, fuel and time wastage, and accidents. These problems directly impact human health, daily life, social life, and psychological states, including lack of control, fatigue, time pressure, and stress. For effective online monitoring and prediction, smart cameras or sensors must be installed at crowded intersections, and path-planning user applications should guide individuals away from traffic jams. The traffic counts or flow hold statistical significance in data analytics, allowing predictions of increasing vehicle numbers over time and planning for traffic infrastructure accordingly. Socioeconomic indicators and policies also play roles in traffic management. The policies based on speed or flow can be implemented to prevent road crashes, motivating us to address the problem of developing accurate and robust traffic prediction methods. In this paper, we are addressing traffic prediction using a deep learning method.
1.1. Related Work
@@
In traffic forecasting, both parametric and non-parametric models are fundamental approaches for predicting future states. The choice of model depends on the size and complexity of the data, as well as the available computational resources. Parametric models include time series analysis, parametric regression, and Kalman filter models. Time series analysis models, such as the auto-regressive integrated moving average (ARIMA), seasonal ARIMA (SARIMA), and exponential smoothing. The non-parametric approaches, including K-nearest neighbors, decision trees, and support vector machine (SVM) have been widely used.
@@
The Box-Jenkins ARIMA model, which transforms non-stationary data to stationary data using the differences in the time series, is frequently incorporated in traffic volume prediction (Hamed et al., 1995). Another approach for prediction in (Okutani and Stephanedes, 1984), uses measured state data to forecast traffic flow, minimizing prediction errors based on the assumption of data stationarity (Kalman filtering theory). These parametric methods are interpretable but struggle to model complex and non-stationary traffic data effectively. Non-parametric models utilize distance metrics to determine the nearest neighbor of the current state to the historical observation in feature space (Smith et al., 2002). These methods are rich in data requirements, less interpretable, and computationally complex. Both parametric and non-parametric methods utilize historical observed data but do not incorporate spatial information. Due to recent advancements in deep learning, significant improvements have been made in traffic forecasting applications. The state-of-the-art methods in traffic prediction consider various factors, such as spatial topology construction, spatial dependency, temporal dependency, and external factors. These advanced methods are designed to better handle the complexities of modern traffic data and provide more accurate predictions over both short (refer to horizons of up to one hour) and long term (beyond one hour) (Van Lint and Van Hinsbergen, 2012).
Spatial topology construction and spatial dependency
@@
Determining a graph structure that specifies the connectivity of spatial locations is known as the adjacency matrix. Usually, the non-Euclidean geodesic information of the sensor network is incorporated in the model using graph neural networks (GNNs). If the graph topology is defined based on the distance between the nodes, this representation is known as a static graph. The semantic adjacency matrix is computed using dynamic time warping (DTW) distance by incorporating the time series of nodes (Fang et al., 2021)(Wang et al., 2024a). The binary adjacency matrix represents the presence of the connection between nodes (Guo et al., 2019). The road network is also included in the distance-based graph structure (Li et al., 2018). The connectivity of a graph is determined using a Gaussian thresholded kernel function and utilizes the weighted adjacency matrix (Zheng et al., 2020). Some literature also built an adjacency matrix using point of interest (POI) similarity (Geng et al., 2019), edge-wise graph (Chen et al., 2020), and free-flow reachable matrix (Cui et al., 2019). In an adaptive graph, the structure of the graph is constructed based on the temporal evolution of the data (Huang et al., 2025). For spatial graph modeling based on the dynamic features, using node embeddings, a self-adaptive adjacency matrix is determined in an end-to-end learning mechanism (Wu et al., 2019)(Bai et al., 2020)(Shao et al., 2022).
@@
In spatial dependencies, we dynamically derive the correlation between nodes of features (Inter-node dependencies) or multiple features of the same node (Intra-node dependencies) from the data. To extract the spatial features of traffic flows from the road network matrix, trafficGAN (Zhang et al., 2019) implements deformable convolution (Dai et al., 2017). Deformable convolution is used to model geometric variations. To capture the spatial dependency, mostly convolutional neural network (CNN) (Zhang et al., 2020)(Lin et al., 2019), and GNN (Li et al., 2018)(Cai et al., 2020)(Bai et al., 2020) are employed in prediction applications. Dynamic graph convolutional recurrent network (Li et al., 2023) incorporates the weighted combination of both static and dynamic graph convolution. The self-attention mechanism (Nie et al., 2024) is also used in the modeling of spatial dependency. In the most recent literature on traffic prediction (Lan et al., 2022)(Guo et al., 2019), the combination of attention and GNN is also encouraged to determine the spatial correlation. The meta-graph attention, which is based on graph attention (GAT) is applied to determine the spatial correlations from traffic data  (Pan et al., 2019). Chebyshev spectral convolution (ChebNet) is incorporated to learn spatial dependencies for spectral graph convolution (Yu et al., 2018). To overcome the smoothing issues in GNNs in deeper networks that impact the long-range dependencies, the spatial-temporal graph ordinary differential equation (STGODE) block is proposed in (Fang et al., 2021). STGODE captures the spatial and temporal dynamic correlations based on ODE. Similarly, neural rough differential equations have been incorporated to capture spatio-temporal dynamics (Choi and Park, 2023). A spatio-temporal transformer graph is used to capture long-term dependencies (Huang et al., 2025).
@@
Traffic prediction involves time series data used to capture sequential relationships. Recurrent neural networks (RNNs) and their optimized variants, such as long short-term memory (LSTM) and gated recurrent units (GRUs), are commonly employed. These models are capable of capturing the long-range temporal dependencies (Li et al., 2023)(Zhang et al., 2022). However, combining dynamic graphs with RNNs makes the model computationally expensive due to a large number of trainable parameters. Temporal convolution filters are also used to extract the temporal information from the data. The attention mechanism (Vaswani et al., 2017) in the spatiotemporal (ST) block is commonly used in predictive models (Guo et al., 2019)(Lan et al., 2022). A combination of RNNs along with the attention mechanism is also employed. GRUs handle the short-term dependencies, while the self-attention mechanism models the long-term dependencies (Shao et al., 2022). Three different types of attention– temporal, spatial, and channel– are applied to capture spatiotemporal correlations (Nie et al., 2024). However, this proposed mechanism increases the computational, time complexity, and number of parameters compared to CNNs. CNNs and temporal convolutional networks (TCNs) have proven more efficient in capturing complex spatiotemporal relationships. The transformer encoder-decoder architecture is implemented for the temporal pattern in (Cai et al., 2020)(Pan et al., 2019). Typically, the encoder is implemented to learn historical traffic information (Han et al., 2024) and the decoder is incorporated to make predictions.
@@
Traffic data consists of complex dynamics that depend on various factors such as socio-economic, land uses land covers, population density, road capacity, and cultural or public events (Li et al., 2022). To apply self-attention to the point of interest (POI) of each region, POI-MetaBlock is included in a deep neural network (DNN) (Wang et al., 2023). To effectively capture temporal correlation, the external weather data is incorporated in DNNs for traffic flow forecasting in (Hou et al., 2021). The non-traffic data, such as rainfall (Jia et al., 2017) is also included in the model. Information such as holidays, weather, and nearby POIs is served as the input to the Large Language Models (LLMs) (Guo et al., 2024). These factors serve as the features of DNNs that assist the network in extracting the spatial and temporal correlation from the traffic data.
@@
Fourier domain analysis is also used in traffic forecasting, where the traffic data is segregated into frequency components. However, the limitation of the Fourier Transform is that the data should be stationary. Sun et al. (Sun et al., 2024) proposed a wavelet learning based decomposition for denoising and determining the frequency components. However, this method may suffer from modes overlapping in high-dynamical systems. So, many signal processing-based methods have been introduced for the decoupling of signals (Wang et al., 2024b). Various methods are commonly used for signal decomposition, including the well-known methods Empirical Mode Decomposition (EMD) (Huang et al., 1998), Ensemble EMD (EEMD) (Wu and Huang, 2009), Complete Ensemble EMD with Adaptive Noise (CEEMDAN) (Torres et al., 2011), Multivariate EMD (MEMD) (Rehman and Mandic, 2010), and Variational Mode Decomposition (VMD) (Dragomiretskiy and Zosso, 2013). The traffic data is decomposed into periodic, residual, and volatile components using the EMD method. A Bi-directional LSTM model is incorporated to model the volatile component (Huang et al., 2022). A hybrid model, combining Improved CCEMDAN (ICCEMDAN) and GRU, is proposed to predict the parking demand (Li and Zhong, 2023)(Tian et al., 2025) used the permutation entropy algorithm to rank intrinsic mode functions (IMFs), derived from ICEEMDAN, IMFs having a small entropy value fed to ARIMA for traffic prediction. Short-term traffic flow is predicted using VMD and LSTM (Lu, 2023). The encoder-decoder architecture is incorporated for feature extraction from mode functions decomposed by VMD in short-term traffic flow prediction (Guo et al., 2023). However, these methods do not incorporate the spatial and topological information. Cross-channel dependencies between decomposed modes are determined by using convolution (Zhang et al., 2025)(Su et al., 2024). These frameworks do not incorporate the graph structure. Existing decomposition methods, such as EMD, EEMD, and CEEMDAN-based approaches, suffer from mode-mixing, while VMD overcomes these challenges by ensuring better frequency separation and adaptability and therefore, offers non-recurrent frequency modes and improved forecasting performance (Zhao et al., 2023).
@@
Figure 1. A visualization showcasing the temporal normalized loss variations on SD region for (a) Horizon 3 (b) Horizon 6 (c) Horizon 12 using our proposed model. Please use Adobe Reader for enhanced and optimal visualization.
1.2. Contributions
To bridge the gap in existing approaches that either ignore the multi-frequency structure of traffic time series or require manual tuning of decomposition parameters, we introduce the Variational Mode Graph Convolutional Network (VMGCN), the first end-to-end framework, that decomposes each node signal into optimally selected variational modes and jointly learns spatiotemporal interactions across these modes using attention-enhanced graph convolutions. In contrast to attention-based spatiotemporal graph convolutional networks (ASTGCN) and Graph WaveNet (GWNET), which learn patterns directly from raw traffic flows, VMGCN leverages frequency-domain insights to disentangle non-stationary effects before applying graph learning. The main contributions of this work are summarized below:
  • In our hybrid framework for spatiotemporal forecasting, we propose to decompose the spatiotemporal traffic data into modes representing the granularity of the data using VMD. The spatial and temporal attention are applied to the decomposed features, followed by ChebNet graph convolution and time convolution to predict the future states over the long term. Such an attention-augmented GCN processes each variational mode separately before fusing both frequency-specific and cross-mode dependencies.
  • We analyse the orthogonality, frequency overlapping, and residual components of decomposed modes of the graph network to reveal the significance of mode components.
  • We also conduct an in-depth analysis of our proposed architecture and formulate a method to determine the best-fit value of the number of modes. Additionally, we highlight the impact assessment of different hyper-parameters on the long-term prediction performance of the model.
  • We evaluate the predication capability of the proposed architecture on the LargeST (Liu et al., 2024) dataset across three different regions using different metrics and demonstrated superior performance as compared to state-of-the-art GNNs. The collective nodes of the LargeST dataset are $6902$, sampled at $15$ minute time interval over one year (containing $241$ million spatiotemporal observations). As an illustration, Fig. 1 demonstrates the prediction loss of our proposed model on the SD region. We also analyze the computational complexity of the proposed method.
The rest of the paper is organized as follows. Section 2 describes the adopted notation and formulates the problem statement. Section 3 presents the detailed architecture of the network used in our framework along with the mathematical formulation of the decomposition method. Section 4 presents the experimental analysis and the improvements enabled by our methods before Section 5 concludes the paper.
2. Preliminaries
2.1. Road Network
@@
In graph theory, a road is modeled by considering different nodes at distinct joints, vertices, or intersections. Each node of the road is connected to another segment or node. The interconnectivity and distance among nodes are incorporated in a graph referred to as a road network, defined as $\mathcal{G}=(V,\,A,\,Y)$, where $V$ is a set of entities or nodes (vertices), $Y$ represents the features, and $A\in\mathbb{R}^{N\times N}$ is the weighted adjacency matrix. Here, $N$ represents the total number of nodes in a traffic prediction application, which could be either the number of sensors or vertices of a road network.
2.2. Temporal Traffic Network Graph
@@
Traffic flow, the number of vehicles per unit of time in a road segment, is a key measure of traffic congestion, which also depends on the width or number of road lanes. A temporal dynamic graph is an arrangement of graph snapshots [$\mathcal{G}^{(1)},\,\mathcal{G}^{(2)},\dots,\,\mathcal{G}^{(T_{w})}$] where $\mathcal{G}^{(t)}=(V^{(t)},\,A^{(t)},\,Y^{(t)})$ includes vertices $V^{(t)}$, an adjacency matrix $A^{(t)}$, and a feature matrix $Y^{(t)}$ (Kazemi, 2022). Here, $T_{w}$ is the time window depending on the historical observation of the signal. The feature matrix $Y^{(t)}$ may include traffic information and road network characteristics. This structure, known as the temporal traffic network graph (TTNG), explains the traffic flow among different nodes from time $[1,T_{w}+N_{H}]$. As an illustration, Fig. 2 depicts the sequence of temporal graphs for a traffic network.
@@
Figure 2. Left side illustrates the spatial distribution of nodes in the San Diego region, as represented in the LargeST dataset. On the right side, the TTNG structure is depicted, where the historical and current input graphs are sampled at time intervals $t=\{1,2,\dots,T_{w}\}$. The output graphs are then predicted for $t=\{T_{w}+1,T_{w}+2,\dots,T_{w}+N_{H}\}$.
2.3. Problem Formulation